The quantity demanded of Good Z depends upon the price of Z (Pz), monthly income (Y), and the price of a related Good W (Pw). Demand for Good Z (Qz) is given by equation 1 below: Qz = 150 - 8Pz + 2Y - 15Pw

Find the demand equation for Good Z in terms of the price for Z (Pz), when Y is $50 and Pw = $6
1
Expert's answer
2020-11-10T10:41:53-0500

Solution:

To get the demand equation for Good Z, substitute the two values into the demand equation given:

"Qz = 150+8Pz+2Y-15Pw"


"Qz = 150+8Pz+(2\times 50) -(15\times 6)"


"Qz = 150+8Pz+100 -90"


"Qz = 150+100-90-8Pz"


"Qz = 160-8Pz"


The demand equation = "Qz = 160-8Pz"